The time (t) of travel of an automobile trip varies inversely as the speed (v).
At an average speed of 65mph, a trip took 4h. The return trip took 5h.
Find the average speed of the return trip.

Answer by ankor@dixie-net.com(22740)

The time (t) of travel of an automobile trip varies inversely as the speed (v).
At an average speed of 65mph, a trip took 4h. The return trip took 5h.
Find the average speed of the return trip.
:
Let s = the speed on the return trip
;
the distance there and back is the same
:
Write a distance equation; dist = speed * time
:
5s = 4(65)
5s = 260
s = 260/5
s = 52 mph on the return trip
:
:
check solution by finding the distances, should be the same
65 * 4 = 260
52 * 5 = 260